# Coverage Axis++: Efficient Inner Point Selection for 3D Shape Skeletonization

### v5

@@ 5. Experimental Results
Dog
@@
Hand-1
$100$ $9.672\%$ $1194$ $4.110\%$ $44$ $3.441\%$ $44$ ${\color[rgb]{0.94140625,0.48046875,0.24609375}\bm{2.328\%}}$
Kitten
$100$ $8.724\%$ $1194$ $5.332\%$ $50$ $3.450\%$ $50$ ${\color[rgb]{0.94140625,0.48046875,0.24609375}\bm{2.962\%}}$
Snake
@@
Average
- $8.003\%$ $4.815\%$ $2.515\%$ ${\color[rgb]{0.94140625,0.48046875,0.24609375}\bm{2.404\%}}$
6. Ablation Study
6.1. Skeletal Point Number
In Sec.4.6, we highlight Coverage Axis++’s practical benefit of flexible specification for the target number of skeletal points. Fig.6 presents our experiments with different target skeletal numbers using three models and three choices of $|V|$. Larger values of $|V|$ result in a more accurate skeletal representation, as observed in the experiments where reconstruction error consistently decreases with increasing $|V|$. However, choosing excessively large $|V|$ contradicts our objective of generating a simple and compact skeleton for shape abstraction. Our results show that specifying a smaller number of skeletal points (e.g., $30$ or $50$) achieves simpler shape abstraction with only a slight increase in reconstruction error. The flexibility allows users to tailor the method for either a simple or more accurate model, depending on the requirement for shape abstraction.
@@ 6.1. Skeletal Point Number
Figure 6. Influence of target numbers $|V|$. We test our model with the target numbers being $30$, $50$ and $70$.
6.2. Dilation Factor
@@
Figure 7. Skeletonization results using different dilation offsets.
6.3. Running Time Analysis
@@
The overall complexity of Algorithm 1 is $O(|S|\cdot|P|\cdot|V|)$. Next, we validate the complexity using different set sizes $|P|$, $|S|$ and $|V|$ on the running time of our method. As shown in Fig. LABEL:fig:time_ablation, Coverage Axis++ has polynomial complexity relationships among various variables, maintaining overall low computational costs.
7. Discussion and Conclusion
@@
In this paper, we present Coverage Axis++, an efficient shape skeletonization method. Unlike existing methods that depend on input watertightness or suffer from high computational costs, Coverage Axis++ addresses these challenges by employing a heuristic algorithm considering both shape coverage and centrality for selecting skeletal points, which ensures an efficient approximation of the MAT while significantly saving the computational cost. Compared with existing methods, our simple yet effective approach produces a compact representation of MAT, offers versatility in skeletonizing various shape representations, allows customizable specification of skeletal points, and achieves highly efficient computation with improved reconstruction accuracy. Extensive experiments across diverse 3D shapes demonstrate the efficiency and effectiveness of Coverage Axis++.
Appendix A Appendix
@@
This supplementary material covers: more experimental results (Sec. A.1), coverage rate of Coverage Axis++ (Sec. A.2), computation complexity analysis on Coverage Axis (Sec. A.3) as well as limitations and future works (Sec. A.4).
A.1. More Experimental Results
@@
As discussed in Sec. 4.6, the proposed method has polynomial complexity with respect to the number of surface samples, denoted as $|S|$. In contrast, the Coverage Axis algorithm showcases exponential complexity in $|S|$. This is evidenced by Tab. 1 of the main paper, where we report the running times of our method and Coverage Axis for various models when $|S|$ is set to $1500$. Furthermore, it can be expected that when increasing the number of surface samples, the performance gap between our method and the Coverage Axis will widen further. We conducted additional experiments by increasing the number of surface samples to $5000$ and comparing the performance of our method with the Coverage Axis across different shapes. The results are summarized in Tab. A5.
@@
A.2. Coverage Rate
In the main paper, we conducted extensive experiments for Coverage Axis++ on various models to demonstrate that our method is capable of extracting compact underlying shape representation of MAT while achieving low reconstruction error, as summarized in Tab. 1. Since the selected skeletal points form a good representation of the medial axis, the corresponding dilated balls should cover a significant portion of the surface samples to build the correspondence between the medial balls and the shape volume. In other words, the coverage rate of our method should be high. The idea has been verified by Coverage Axis, as it directly requires a $100\%$ coverage rate by imposing a hard constraint during the optimization. We record the coverage rates of each model for Coverage Axis++, which are summarized in Tab. A4. The result shows that our method consistently achieves very high coverage rates, with an average of $96.134\%$.
Table A4. Coverage rate of Coverage Axis++. The inputs are meshes. $|V|$: The number of skeletal points. $\overleftrightarrow{\epsilon}$: Two-sided HD between original surface and reconstruction.
Model
$|V|$
$\overleftrightarrow{\epsilon}$ Coverage rate
Bear-1
$68$
$4.047\%$ $98.662\%$
Bear-2
$49$
$4.572\%$ $98.461\%$
Bird
$83$
$2.560\%$ $95.914\%$
Bug
$137$
$4.511\%$ $98.274\%$
Bunny
$73$
$5.771\%$ $97.650\%$
Camel
$93$
$5.556\%$ $96.919\%$
Chair
$115$
$2.947\%$ $96.082\%$
Crab-1
$93$
$2.744\%$ $97.874\%$
Crab-2
$107$
$2.825\%$ $98.608\%$
Cup
$243$
$9.247\%$ $95.411\%$
Dinosaur
$66$
$2.672\%$ $99.331\%$
Elephant
$94$
$3.874\%$ $98.079\%$
Fandisk
$127$
$4.067\%$ $92.029\%$
Femur
$26$
$3.095\%$ $97.536\%$
Fish
$43$
$4.809\%$ $88.587\%$
Giraffe
$71$
$4.208\%$ $97.335\%$
Guitar
$71$
$2.497\%$ $96.804\%$
Hand
$47$
$1.975\%$ $95.143\%$
Human
$46$
$2.105\%$ $98.869\%$
Lifebuoy
$33$
$3.324\%$ $96.151\%$
Octopus
$74$
$3.005\%$ $97.989\%$
Plane
$44$
$4.306\%$ $83.389\%$
Pot
$90$
$5.278\%$ $95.786\%$
Rocker
$112$
$3.652\%$ $96.944\%$
Seahorse
$60$
$5.168\%$ $96.144\%$
Spectacles
$97$
$3.080\%$ $96.075\%$
Spider
$54$
$4.493\%$ $95.987\%$
Vase
$118$
$3.875\%$ $95.710\%$
Average
$3.938\%$ $96.134\%$
A.3. Complexity Analysis on Coverage Axis

### v7

@@ 5.1.1 Comparison with Coverage Axis
Bird $83$ $2.3$ $2.361\%$ $3.936\%$ $3.936\%$ 3.1 $2.519\%$ $4.034\%$ $4.034\%$ $2.3$ $2.306\%$ $2.560\%$ $2.560\%$
Bunny $73$ $16.6$ $3.474\%$ $4.686\%$ $4.686\%$ 19.7 $3.501\%$ $3.762\%$ $3.762\%$ $1.5$ $4.860\%$ $5.771\%$ $5.771\%$
Camel $93$ $2.8$ $2.204\%$ $2.468\%$ $2.468\%$ 4.1 $2.047\%$ $2.110\%$ $2.110\%$ $2.5$ $2.175\%$ $5.556\%$ $5.556\%$
Dinosaur $66$ $8.1$ $1.930\%$ $2.444\%$ $2.444\%$ 9.9 $1.942\%$ $3.408\%$ $3.408\%$ $1.4$ $2.308\%$ $2.672\%$ $2.672\%$
Dog $69$ $1.3$ $2.299\%$ $3.781\%$ $3.781\%$ 1.7 $2.173\%$ $2.725\%$ $2.725\%$ $1.5$ $3.353\%$ $3.799\%$ $3.799\%$
Fandisk $127$ $2.9$ $4.888\%$ $5.840\%$ $5.840\%$ 3.8 $4.911\%$ $5.920\%$ $5.920\%$ $4.6$ $3.849\%$ $4.067\%$ $4.067\%$
Hand $47$ $21.9$ $1.441\%$ $1.087\%$ $1.441\%$ 25.5 $2.173\%$ $2.208\%$ $2.208\%$ $1.1$ $1.974\%$ $1.975\%$ $1.975\%$
Horse $81$ $22.9$ $1.973\%$ $3.683\%$ $3.683\%$ 24.4 $2.362\%$ $3.853\%$ $3.853\%$ $1.9$ $2.382\%$ $2.658\%$ $2.658\%$
Human 46 $1.6$ $2.006\%$ $2.155\%$ $2.155\%$ 2.1 $1.545\%$ $2.217\%$ $2.217\%$ $0.9$ $2.062\%$ $2.105\%$ $2.105\%$
Neptune $53$ $5.0$ $2.379\%$ $11.372\%$ $11.372\%$ 7.2 $4.229\%$ $7.146\%$ $7.146\%$ $1.2$ $3.227\%$ $9.989\%$ $9.989\%$
Seahorse $60$ $7.6$ $1.908\%$ $4.731\%$ $4.731\%$ 8.9 $1.806\%$ $4.557\%$ $4.557\%$ $1.4$ $2.296\%$ $5.168\%$ $5.168\%$
Vase $118$ $6.2$ $2.826\%$ $4.839\%$ $4.839\%$ 8.0 $3.038\%$ $3.893\%$ $3.893\%$ $3.3$ $3.575\%$ $3.875\%$ $3.875\%$
5.1.2 Comparison with MATFP
We further compare our method with the recent work MATFP [WWWG22]. MATFP aims to compute MAT with feature preservation, and it only takes watertight mesh input. Besides, similar to the Coverage Axis, the user cannot explicitly specify the final number of skeletal points. As shown in Fig. 3, we find MATFP typically yields results with redundant points. In comparison, Coverage Axis++ produces more compact results while using less computation time (13.11s v.s. ours 2.61s). Since MATFP is a multi-stage approach, we detail the running time of each stage in Appendix B.
5.1.3 Comparison with Q-MAT and Q-MAT+
We compare our method with two representative methods for MAT computation given the mesh inputs: Q-MAT [LWS15] and its variant Q-MAT+ [PWG19]. Both two methods take watertight meshes as inputs and cannot handle point clouds or polygon soups. Tab. 2 shows that Coverage Axis++ performs comparably or superior to Q-MAT and Q-MAT+ in terms of reconstruction accuracy and running time. Note that both Q-MAT and Q-MAT+ require a watertight mesh for Voronoi initialization, while our method supports randomly generated candidate points inside the volume.
5.1.4 Comparison with SAT
SAT [MGP10] encounters a major drawback as it tends to favor a dense representation with numerous vertices, hindering the generation of a simple and compact skeleton of the given shape. Note SAT only excels specifically with watertight surface meshes. We test with different scaling factors, $\alpha=1.1$, $1.5$, and $2.0$, revealing that with a smaller $\alpha$, SAT achieves high accuracy but results in a representation with a large number of vertices (Fig.3(a-c)). Using a larger $\alpha$ for higher abstraction compromises shape structure and introduces significant approximation errors. As a comparison, Coverage Axis++ achieves a balance between fidelity, efficiency, and compression abilities.
5.2 Point Cloud Inputs
5.2.1 Comparison with Neural Skeleton
We compare Coverage Axis++ with Neural Skeleton [CD23] given the point cloud inputs. Neural Skeleton first trains a SDF w.r.t. the input point cloud, then samples uniform surface points and skeletal points from the implicit field. Finally, the skeleton mesh is recovered using a cover-set formulation [DLX22] solved as a Mixed-Integer Linear Program. The implicit field improves the robustness of the Neural Skeleton in handling challenging inputs. As shown in Tab. 3, our method produces results with higher accuracy. Moreover, the average running time of Coverage Axis++ and NeuralSkeleton is 1.63s v.s. 59.27s. The detailed runtime statistics of Neural Skeleton are given in Appendix B. Note that the Neural Skeleton does not support specifications for the number of skeletal points.
5.2.2 Comparison with Point2Skeleton
Compared with Point2Skeleton [LLL21], a learning-based skeletonization method for point clouds, our approach exhibits better performance. As shown in Fig. 5 and Tab. 3, Point2Skeleton, trained on a large dataset ShapeNet [CFG15], falls short on shapes unseen during training and struggles with higher topology complexity. However, our method demonstrates robustness, generating high-quality skeletal representations with accurate geometries and faithful structures across diverse shapes.
5.2.3 Comparison with Deep Point Consolidation
As shown in Fig. 5 and Tab. 3, Deep Point Consolidation (DPC) [WHG15] can only produce unstructured inner points without connections and the reconstruction results exhibit large errors. Even worse, the reconstructed topologies by DPC are usually inconsistent with the original input. In contrast, our reconstruction results reach better approximation accuracy with respect to the original geometry using fewer skeletal points.
Table 3: Comparison on shape approximation errors among Point2Skeleton (P2S), Deep Point Consolidation (DPC), Neural Skeleton, Coverage Axis and Coverage Axis++.
Model P2S DPC Neural Skeleton Coverage Axis Coverage Axis++
$|V|$ $\overleftrightarrow{\epsilon}$ $|V|$ $\overleftrightarrow{\epsilon}$ $|V|$ $\overleftrightarrow{\epsilon}$ $|V|$ $\overleftrightarrow{\epsilon}$ $|V|$ $\overleftrightarrow{\epsilon}$
Ant-2 $100$ $16.412\%$ $1194$ $8.863\%$ 54 $3.993\%$ $58$ $2.350\%$ $58$ ${\color[rgb]{240,123,63}\bm{2.113\%}}$
Bottle $100$ $2.955\%$ $1194$ $2.752\%$ 13 $2.390\%$ $14$ $2.956\%$ $14$ $2.599\%$
@@ 5.2.3 Comparison with Deep Point Consolidation
Dolphin $100$ $5.925\%$ $1194$ $7.916\%$ 36 $3.269\%$ $49$ $1.971\%$ $49$ ${\color[rgb]{240,123,63}\bm{1.876\%}}$
Fertility $100$ $8.162\%$ $1194$ $4.226\%$ 76 $5.909\%$ $79$ $3.428\%$ $79$ ${\color[rgb]{240,123,63}\bm{3.102\%}}$
@@
Hand $100$ $9.672\%$ $1194$ $4.110\%$ 39 $4.819\%$ $44$ $3.441\%$ $44$ ${\color[rgb]{240,123,63}\bm{2.328\%}}$
Kitten $100$ $8.724\%$ $1194$ $5.332\%$ 41 $3.883\%$ $50$ $3.450\%$ $50$ ${\color[rgb]{240,123,63}\bm{2.962\%}}$
@@
Average - $8.003\%$ $4.815\%$ $3.528\%$ $2.515\%$ ${\color[rgb]{240,123,63}\bm{2.404\%}}$
6 More Discussions
[…]
Figure 6: Influence of target numbers $|V|$. We test our model with the target numbers being $30$, $50$ and $70$.
6.1 Skeletal Point Number
In Sec. 4.7.2, we highlight Coverage Axis++’s practical benefit of flexible specification for the target number of skeletal points. Fig. 6 presents our experiments with different target skeletal numbers using three models and three choices of $|V|$. Larger values of $|V|$ result in more accurate skeletal representations, as observed in the experiments where reconstruction error consistently decreases with increasing $|V|$. However, choosing excessively large $|V|$ contradicts our objective of generating a simple and compact skeleton for shape abstraction. Our results show that specifying a smaller number of skeletal points (e.g., $30$ or $50$) achieves simpler shape abstraction with only a slight increase in reconstruction error. The flexibility allows users to tailor the method for either a simple or more accurate model, depending on the requirement for shape abstraction.
6.2 Dilation Factor
@@
Figure 7: Skeletonization results using different dilation offsets.
[…]
Figure 8: We conducted ablation experiments on different weights. We test our method using different weights corresponding to the centrality score, uniformity score, and coverage score, e.g., $(1.0,1.0,0.2)$ refers to $Score_{i}=1.0\cdot Cov_{i}+1.0\cdot Unif_{i}+0.2\cdot Centr_{i}$. We visualize the skeletal results and evaluate the reconstruction error. We test with Human and Crab models.
[…]
Figure 9: Performance of different combinations of the three scores. (a) coverage score. (b) coverage + uniformity score. (c) coverage + uniformity + centrality score (Eq. 9).
6.3 Ablation on Scoring Terms
We evaluate our method with different scoring terms defined as different combination schemes of the coverage (Eq. 5), uniformity (Eq. 6)and centrality (Eq. 7) scores on the Crab and Human models, where we set $|V|=60$ for both models. From Fig. 9, we observe that the inclusion of the uniformity score results in skeletonizations with fewer sharp triangles, indicating a more uniform distribution of selected points. Furthermore, incorporating the centrality score enables the selected skeletal points to capture more accurate MAT information, leading to a more compact and regular skeletonization. Consequently, incorporating all three scores, as shown in Eq. 9 yields improved skeletonization outcomes, i.e., compact and high-accuracy shape approximation.
Although using equal weights for the terms in Eq.9 effectively achieves compact skeletonization and high reconstruction accuracy in our main experiments without requiring special weight adjustments, we investigate the impact of different weighting schemes for each term. Concretely, we scale the weight of each term down and up by a factor of 5, while keeping the weights of the other terms unchanged, to investigate the influence of different terms. We test our algorithm using two models: Crab and Human. As shown in Fig. 8, we observe that applying a moderate uniformity weight significantly enhances the spatial distribution uniformity of the model vertices. However, using excessively large centrality score weights may deteriorate the regular structure of skeletal points, resulting in increased reconstruction error.
6.4 Shape Skeletonization for Triangle Soups
We evaluate our method for shape skeletonization using triangle soup inputs. For the experiment, we apply a 0.25% perturbation relative to the bounding box to each vertex of the triangles. The skeletonization process is identical to that used for handling point cloud inputs. In this experiment, we sample $1500$ surface points from the soup and use $10000$ randomly generated candidate points inside the volume, as we described in Sec. 4.5. As shown in Fig. 10, our method yields high-quality skeletonization and shape approximation results with the triangle soup inputs.
[…]
Figure 10: Shape skeletonization results for triangle soup. (a) Input triangle soup. (b) Skeletonization result. (c) Shape reconstruction.
6.5 Running Time Analysis
@@
The overall complexity of Algorithm 1 is $O(|S|\cdot|P|\cdot|V|)$. Next, we validate the practical running time complexity of our method using different set sizes $|P|$, $|S|$, and $|V|$. As shown in Fig. LABEL:fig:time_ablation, Coverage Axis++ has polynomial complexity relationships among various variables, maintaining overall low computational costs.
6.6 Limitations and Future Works
Coverage Axis++ approximates the whole shape by an abstraction of local geometry with the union of the local geometries forming the entire shape. One limitation of Coverage Axis++ is that its whole process is conducted without a hard constraint forcing the consistency between the selected medial balls and the shape surface (See coverage rate statistics in Appendix. C), like that in Coverage Axis [DLX22]. Although local approximation for shape skeletonization has been evidenced by some previous works [LLL21, GYY23, ACK01], we acknowledge that the inclusion of a global constraint imposes a much stricter guarantee for the shape correspondence.
Since the goal of this paper is to compute an overall compact skeletonization result, our method may encounter challenges when handling shapes like CAD models that have extremely thin structures or sharp on-surface features (see the inset figure). In the future, introducing additional terms to preserve sharp features during shape skeletonization would be promising. For example, an algorithm can be developed to build up the connection of sharp features and the medial axis, ensuring the skeletonization result retains the sharp features on the surface. Our method can generate a compact representation in terms of shape coverage using medial balls as primitives. We admit that while minimizing the redundancy of skeletal points is highly desirable, there is still potential to further reduce the number of skeletal points. For example, planar components such as the body of a guitar, the seat of a chair, and the base of a fertility statue are currently oversampled because the coverage capability is limited by the medial balls that are isotropic, as shown in Fig. Coverage Axis++: Efficient Inner Point Selection for 3D Shape Skeletonization. Future work could explore the use of different types of bounding volume primitives, such as slab meshes [LWS15], ellipsoids [LCWK07], or medial skeletal diagrams [GWM23] instead of medial balls, to achieve a more compact representation with fewer skeletal points, while maintaining the coverage optimization goal.
Next, we discuss more potential applications and future works. In this paper, we present an effective and efficient method for skeletal point selection and MAT computation. The crucial geometrical and structural information encoded in the MAT could be helpful in various downstream applications. For instance, in the task of shape segmentation [LLL20], the MAT can be leveraged for identifying various types of junctions between different parts of a 3D shape using information like manifold branch and dimensional change of the MAT without relying on 3D ground truth supervision for segmentation. In the future, we believe that a part-level shape prior derived from the MAT through self-supervised learning could benefit part-aware shape generation and reconstruction [LLL24]. Meanwhile, the discriminative feature represented by the MAT could also facilitate the 3D shape recognition as shown by [HWQ19]. MAT has demonstrated its effectiveness in deep, topology-aware generation for 3D shapes [PGT24] with the neural skeleton-based representation encoding information on both shape topology and geometry. It has also been evidenced that the topological and geometrical information of the medial mesh can be useful in remeshing and volumetric subdivision, e.g., achieving all-hexahedral mesh generation of the given shape, making the method more general than previous methods based on curve-skeleton [ZXW24].
Even the point selection part of our method could serve as an effective point sampling strategy for compact shape encoding in wide applications. For example, the point selection method [DLX22] has been shown to be effective in enhancing computational and memory efficiency for 3D shape representation, e.g., TetSphere Splatting [GWHM24]. More potential applications of our point selection method can be those point-based methods. Specifically, one may achieve compact 2D/3D Gaussian Splatting [HYC24, KKLD23], also known as Gaussian pruning, given a properly defined coverage score during the training. Actually, our point selection algorithm can serve as a general sampling approach applicable to various tasks. As long as an original input representation can be expressed as candidate points and an associated covering matrix are properly defined, thanks to the SCP, our method could yield a more compact representation for the given redundant input.
7 Conclusion
@@
In this paper, we present Coverage Axis++, an efficient shape skeletonization method. Unlike existing methods that depend on input watertightness or suffer from high computational costs, Coverage Axis++ addresses these challenges by employing a heuristic algorithm considering shape coverage, uniformity, and centrality for selecting skeletal points, which ensures an efficient approximation of the MAT while significantly saving the computational cost. Compared with existing methods, our simple yet effective approach produces a compact representation of MAT, offers versatility in skeletonizing various shape representations, allows customizable specification of skeletal points, and achieves highly efficient computation with improved reconstruction accuracy. Extensive experiments across diverse 3D shapes demonstrate the efficiency and effectiveness of Coverage Axis++.
8 Acknowledgements
The authors would like to thank Baorong Yang, Mengyuan Ge, and Ningna Wang for the fruitful discussion and the anonymous reviewers for their valuable comments and suggestions. Zhiyang Dou is supported by the Innovation and Technology Commission of the HKSAR Government under the InnoHK initiative. Shiqing Xin is supported by the National Natural Science Foundation of China (Ref: 62272277). This research is partly supported by Innovation and Technology Commission (Ref: ITS/319/21FP) and Research Grant Council (Ref: 17210222), Hong Kong.
@@ 8 Acknowledgements
This supplementary material covers: more experimental results (Sec. A), more detailed running time statistics of existing methods (Sec. B), coverage rate of Coverage Axis++ (Sec. C), computation complexity analysis on Coverage Axis (Sec. D), and an additional discussion on the skeletal point connectivity of MAT (Sec. E).
Appendix A More Experimental Results
@@
As discussed in Sec. 4.7.1, the proposed method has polynomial complexity with respect to the number of surface samples, denoted as $|S|$. In contrast, the Coverage Axis algorithm showcases exponential complexity in $|S|$. This is evidenced by Tab. 1 of the main paper, where we report the running times of our method and Coverage Axis for various models when $|S|$ is set to $1500$. Furthermore, it can be expected that when increasing the number of surface samples, the performance gap between our method and the Coverage Axis will widen further. We conducted additional experiments by increasing the number of surface samples to $5000$ and comparing the performance of our method with the Coverage Axis across different shapes. The results are summarized in Tab. D4.
@@
Appendix B More Detailed Running Time Statistics.
We provide more statistics of the running time of Neural Skeleton [CD23], MATFP [WWWG22]. All evaluations are performed on the same machine to ensure a fair comparison.
Table B1: Detailed running time statistics of MATFP [WWWG22] using mesh inputs.
Model Stage1 (s) Stage2 (s) Stage3 (s) Stage4 (s) Total Time (s)
Vase 1.62 3.30 0.66 2.69 8.27
Hand 4.15 5.42 1.57 3.54 14.68
Dog 0.59 0.92 1.51 7.53 10.55
Fertility 2.32 4.12 2.19 8.33 16.96
Crab 2.49 4.90 1.48 6.25 15.12
Average 2.23 3.73 1.48 5.67 13.11
Table B2: Detailed running time statistics of Neural Skeleton [CD23] using point cloud inputs.
Model Network Optimization Time (s) Coverage Skeleton Time (s) Total Time (s)
Ant-2 9.95 10.53 20.48
Bottle 9.88 11.61 21.49
Chair-2 9.86 121.52 131.38
Dog 9.96 15.42 25.38
Dolphin 10.07 24.88 34.95
Fertility 9.92 108.15 118.07
Guitar 9.86 107.32 117.18
Hand-1 10.01 21.36 31.37
Kitten 9.89 46.88 56.77
Snake 9.85 25.74 35.59
Average 9.92 49.39 59.27
Appendix C Coverage Rate
In the main paper, we conducted extensive experiments for Coverage Axis++ on various models to demonstrate that our method is capable of extracting compact underlying shape representation of MAT while achieving low reconstruction error, as summarized in Tab. 1. Since the selected skeletal points form a good representation of the medial axis, the corresponding dilated balls should cover a significant portion of the surface samples to build the correspondence between the medial balls and the shape volume. In other words, the coverage rate of our method should be high. The idea has been verified by Coverage Axis, as it directly requires a $100\%$ coverage rate by imposing a hard constraint during the optimization. We record the coverage rates of each model for Coverage Axis++, which are summarized in Tab. C3. The result shows that our method consistently achieves very high coverage rates, with an average of $96.134\%$.
Table C3: Coverage rate of Coverage Axis++. The inputs are meshes. $|V|$: The number of skeletal points. $\overleftrightarrow{\epsilon}$: Two-sided HD between original surface and reconstruction.
Model $|V|$ $\overleftrightarrow{\epsilon}$ Coverage rate
Bear-1 $68$ $4.047\%$ $98.662\%$
Bear-2 $49$ $4.572\%$ $98.461\%$
Bird $83$ $2.560\%$ $95.914\%$
Bug $137$ $4.511\%$ $98.274\%$
Bunny $73$ $5.771\%$ $97.650\%$
Camel $93$ $5.556\%$ $96.919\%$
Chair $115$ $2.947\%$ $96.082\%$
Crab-1 $93$ $2.744\%$ $97.874\%$
Crab-2 $107$ $2.825\%$ $98.608\%$
Cup $243$ $9.247\%$ $95.411\%$
Dinosaur $66$ $2.672\%$ $99.331\%$
Elephant $94$ $3.874\%$ $98.079\%$
Fandisk $127$ $4.067\%$ $92.029\%$
Femur $26$ $3.095\%$ $97.536\%$
Fish $43$ $4.809\%$ $88.587\%$
Giraffe $71$ $4.208\%$ $97.335\%$
Guitar $71$ $2.497\%$ $96.804\%$
Hand $47$ $1.975\%$ $95.143\%$
Human $46$ $2.105\%$ $98.869\%$
Lifebuoy $33$ $3.324\%$ $96.151\%$
Octopus $74$ $3.005\%$ $97.989\%$
Plane $44$ $4.306\%$ $83.389\%$
Pot $90$ $5.278\%$ $95.786\%$
Rocker $112$ $3.652\%$ $96.944\%$
Seahorse $60$ $5.168\%$ $96.144\%$
Spectacles $97$ $3.080\%$ $96.075\%$
Spider $54$ $4.493\%$ $95.987\%$
Vase $118$ $3.875\%$ $95.710\%$
Average - $3.938\%$ $96.134\%$
Appendix D Complexity Analysis on Coverage Axis